# Stochastic gravitational-wave background search using data from five pulsar timing arrays

### v1

@@ Abstract
HD with free $\gamma_{gw}\qquad$ $14.28^{+0.12}_{-0.12}\qquad$ $3.32^{+0.28}_{-0.28}$
HD with $\gamma_{gw}=13/3\qquad$ $14.72^{+0.05}_{-0.05}\qquad$
Table 2: Median posteriors values for $A_{gw}$ and $\gamma_{gw}$, and their 68% containment intervals (gray bands in Fig. 3).
Happily, these posteriors, summarized in Table 2, are consistent with the individual PTA results [7, 10, 9, 6]. They are more constraining than the posteriors obtained by the individual PTAs: the uncertainty ellipses in the bottom-left plot of Fig. 2 are smaller than those obtained by intersecting the individual PTA posteriors [55]. For example, the $\text{area}=\int\!{\rm{d}}\gamma_{gw}\int\!{\rm{d}}\log_{10}(A_{gw})$ of our 95.5% containment ellipse (for the HD hypothesis) is 45% of that shown in [55, Fig. 1, right]. As explained earlier, this is as expected for an SGWB: joint analysis yields more evidence and is more constraining than naive intersection of the separate PTA results.
To claim an SGWB detection, there are several hurdles to overcome [12]. One must show that the data is consistent with the HD signal hypothesis, and that it is very unlikely (at the $5\sigma$ level) to arise under the CURN hypothesis. Clearly, the estimates of $A_{GW}$ and $\gamma_{GW}$ cannot help with this, since they are almost identical under both hypotheses.
One way to assess significance is through null-hypothesis testing [56] using a frequentist statistic. We use quadratic statistics of the form $D(z)=z^{\mathsf{T}}{\mathbf{Q}}z$, where $z$ is the data column vector, and consider three choices of the square filter matrix $\mathbf{Q}$} (see [27, Eq. (1)]). The optimal statistic (OS) selects $\mathbf{Q}$} to maximize the signal-to-noise ratio (SNR) [28]. The Neyman–Pearson (NP) statistic selects $\mathbf{Q}$} to maximize the detection probability at fixed false-alarm probability $p$. The NP minimum-variance (NPMV) statistic adopts the NP matrix but sets all diagonal entries of $\mathbf{Q}$} to zero, eliminating autocorrelation contributions.
Normally, the NP statistic would be preferred. However, because its filter matrix $\mathbf{Q}$} contains both diagonal (autocorrelation) and off-diagonal (cross-correlation) terms, it is not robust against errors in modeling pulsar noise. In contrast, the OS and NPMV statistics use filter matrices that contain only off-diagonal terms, making them more robust. Details may be found in [27].
@@
The values of these statistics and their false-alarm probabilities $p$ depend on the observational data and on the (inferred) pulsar-noise models under the null (CURN) hypothesis. While the data are fixed, the models are uncertain and vary across posterior samples, so we adopt a noise-marginalized approach [28, 29]. For each posterior-sample CURN universe, we compute the three detection statistics and evaluate their false-alarm probabilities analytically using the appropriate generalized $\chi^{2}$ distribution [33, 32]. We then form a histogram of the resulting $p$ values.
We first discuss the NPMV statistic, whose posterior $p$-value histogram is shown in Fig. 3. The mean and median false-alarm probabilities are $p=3.4\times 10^{-3}$ and $p=7.2\times 10^{-4}$ for free $\gamma_{gw}$ and $p=2.9\times 10^{-3}$ and $p=9.0\times 10^{-4}$ for fixed $\gamma_{gw}=13/3$. Respectively, these correspond to $2.7\sigma$, $3.2\sigma$, $2.8\sigma$, and $3.1\sigma$ equivalent Gaussian significances. The most likely range is shown in Fig. 3 as the band that contains 68% of the posterior points about the median. These correspond to significances in the range from $2.5-3.8\sigma$ for free $\gamma_{gw}$ and $2.6-3.7\sigma$ for $\gamma_{gw}=13/3$.
@@
OS $\sigma-$units: 4.0 4.5 (4.0, 5.1) (3.4, 5.6)
$\log_{10}p$: 4.4 5.5 (4.4, 6.7) (3.5, 8.0)
NP $\sigma-$units: 2.6 3.1 (2.5, 3.7) (1.9, 4.2)
$\log_{10}p$: 2.4 3.0 (2.2, 3.9) (1.5, 4.9)
NPMV $\sigma-$units: 2.7 3.2 (2.5, 3.8) (2.0, 4.5)
$\log_{10}p$: 2.5 3.1 (2.3, 4.2) (1.6, 5.4)
@@
The corresponding histograms for the NP and OS false-alarm probabilities $p$ have a similar appearance, so are not shown. Instead, we summarize the most important information for all three statistics in Table 3. None has a mean false-alarm probability $p$ small enough to exceed the $5\sigma$ threshold for confident detection [12]. However, it is notable that our analysis yields a traditional OS detection significance of $\approx 4.0\sigma$, which is substantially higher than the $\approx 3.2\sigma$ significance reported by NG [34]. This indicates that the “direct data combination” method is working as would be expected in the presence of an SGWB.
Another way to quantify detection confidence is through the Bayes factor (BF), defined as the ratio of the evidence for the HD model to that for the CURN model [12]. The evidence is the likelihood marginalized over the prior [57]. In our case, evaluating this marginalization integral is challenging because the parameter space is high-dimensional: $4\times 121+2=486$ dimensions when $\gamma_{gw}$ is free (or $485$ when $\gamma_{gw}$ is fixed to $13/3$). To compute the evidence, we use the generalized stepping-stone sampling method [58, 59, 60], which progressively “broadens” the sharply peaked likelihood using $K=8$ chains with K1=7K-1=7 steps, while also restricting the integration region in parameter space.
This yields $\ln(\mathrm{BF})=10.5\pm 1.0$, corresponding to $\mathrm{BF}\approx 40{\kern 0.83298pt}000$, which is substantially larger than the $\mathrm{BF}\approx 200$ previously reported by NG [9]. For a one-dimensional Gaussian random variable, this BF corresponds to a significance of $\sqrt{2\ln(\mathrm{BF})}\,\sigma\approx 4.6\sigma$ [29, Eq. (8)], which is below that required for a confident detection.
@@
Another detection confidence test is to estimate the inter-pulsar correlation pattern from the data [12]. The mean correlation $\Gamma_{ab}$ between pulsars $a$ and $b$ arising from an SGWB should follow the HD prediction [25, 26, 2]
@@
Our HD reconstruction adopts the NG [9] choice of 15 angular bins. The $121\times 120/2=7{\kern 0.83298pt}260$ pulsar pairs are divided equally: each bin contains 484 pairs. For each posterior sample, the correlations of the pulsar pairs are weighted as in [62], giving $20{\kern 0.83298pt}480$ mean correlations $\mu$ and their variances $\sigma^{2}_{\mu}$ for each bin. For each bin, we then construct and average $20{\kern 0.83298pt}480$ different Gaussian probability distributions. The resulting 15 (posterior estimators of the) probability distributions of the correlation are sharply peaked. Fig. 4 shows their means (very close to medians) and standard deviations. The reconstruction matches the HD curve: the reduced $\chi^{2}$ of $0.84$ is consistent with normal statistical fluctuations ($p$-value $0.63$ for 15 degrees of freedom).
Conclusion — We have conducted a search for a stochastic gravitational-wave background (SGWB) using data from five PTAs. The analysis is straightforward in the PTA context: we apply only minor adjustments to the pulsar noise models and discard a small fraction of data. In essence, we repeat an established analysis using a larger dataset.
@@
Our results provide the strongest evidence to date that an SGWB contributes to the observed pulsar pulse arrival-time fluctuations. However, the statistical significance of this evidence—assessed using multiple frequentist and Bayesian methods—remains too low to support a definitive detection.
@@
Our use of existing data-analysis pipelines and infrastructures makes it straightforward to compare our results with those of the individual PTAs, but it has drawbacks. For example, these pipelines assume that, for a stationary process, the Fourier amplitudes in different frequency bins are statistically independent. While this is not strictly correct [61, see Conclusion], the resulting effects may be negligible or readily corrected [43]. Similarly, our reconstruction of the HD curve is suboptimal because the implementation of an optimal method [61] is still in progress [44].
@@
As part of its upcoming data release, the IPTA is re-analyzing the five-PTA plus CPTA dataset to produce an IPTA-DR3 release (new .par and .tim files). If the primary outcome of this re-analysis is the adoption of common astrophysical models for the pulsars, then—per the arguments above—we expect an SGWB search of the five-PTA data to yield results consistent with ours. But if the re-analysis reveals that some data are incorrect and must be discarded, the results could differ. In this sense, the reliability of our conclusions depends on the correctness of the five existing PTA analyses.
@@
We have deliberately kept this analysis as straightforward as possible and have not “iterated” it. This departs from standard practice in pulsar astronomy, where models—for example, pulsar timing parameters—are iteratively refit for improvement. Here, we avoid such iterative refinement and tuning of the full PTA search pipeline. Past experience with weak-signal detection in the LIGO context, as well as theoretical studies in the PTA context [11], suggests that such iteration can be misleading.
@@
Because they are not public, we have not been able to include CPTA data [8] in our analysis. The CPTA data come from the Five-hundred-meter Aperture Spherical Telescope (FAST) in Guizhou Province, China. FAST is the world’s most sensitive radio telescope, and its data are of very high quality. Including them in future analyses would be both interesting and informative.
Note: the complete code used for this work will be posted publicly at the time of publication.
@@
Acknowledgments – We thank Rutger van Haasteren for many useful conversations and discussions, for his encouragement with this project, and for his comments on various drafts of the manuscript. We also thank Stanislav Babak, Siyuan Chen, Colin Clark, Oliver Dodge, Boris Goncharov, Michael Kramer, Joseph Romano, and Michele Vallisneri for helpful feedback.

### v2

@@ Abstract
HD with free $\gamma_{gw}\qquad$ $14.45^{+0.13}_{-0.15}\qquad$ $3.76^{+0.31}_{-0.29}$
HD with $\gamma_{gw}=13/3\qquad$ $14.72^{+0.04}_{-0.04}\qquad$
Table 2: Median posterior values for $A_{gw}$ and $\gamma_{gw}$, and their 68% containment intervals (gray bands in Fig. 3).
These posteriors, summarized in Table 2, are consistent with the individual PTA results [7, 10, 9, 6]. They are more constraining than the posteriors obtained by the individual PTAs: the uncertainty ellipses in the bottom-left plot of Fig. 2 are smaller (by about a factor of two) than those obtained by intersecting the individual PTA posteriors [57]. For example, the $\text{area}\equiv\int\!{\rm{d}}\gamma_{gw}\int\!{\rm{d}}\log_{10}(A_{gw})$ of our 95.5% containment ellipse for the HD hypothesis is 48% of that shown in [57, Fig. 1, right]. As explained earlier, this is as expected for an SGWB: joint analysis yields more evidence and is more constraining than naive intersection of the separate PTA results.
The PTA community has established basic criteria for claiming a detection [12]. Among these, one must show that the data is consistent with the HD signal hypothesis, and very unlikely (at the $5\sigma$ level) to arise under the CURN hypothesis. The estimates of $A_{GW}$ and $\gamma_{GW}$ cannot help with this, since they are very similar under both hypotheses.
One way to assess significance is the frequentist approach: null-hypothesis testing [58] using a statistic. This computes the (noise-marginalized false-alarm) probability $p$ that signal-like data would occur under the null hypothesis. We use quadratic statistics of the form $D(z)=z^{\mathsf{T}}{\mathbf{Q}}z$, where $z$ is the data column vector, and consider three choices of the square filter matrix $\mathbf{Q}$} (see [28, Eq. (1)]). The optimal statistic (OS) selects $\mathbf{Q}$} to maximize the signal-to-noise ratio (SNR) [30, 33, 28]. The Neyman–Pearson (NP) statistic selects $\mathbf{Q}$} to maximize the detection probability at fixed false-alarm probability $p$ [28]. The NP minimum-variance (NPMV) statistic [28] adopts the NP matrix but sets all diagonal entries of $\mathbf{Q}$} to zero, eliminating autocorrelation contributions.
Why? Normally, the NP statistic would be preferred. However, because its filter matrix $\mathbf{Q}$} contains both diagonal (autocorrelation) and off-diagonal (cross-correlation) terms, it is not robust against errors in modeling pulsar noise. In contrast, the OS and NPMV statistics use filter matrices that contain only off-diagonal terms, making them more robust. Details may be found in [28].
@@
The detection statistic values and their false-alarm probabilities $p$ depend on the observational data and (inferred) pulsar-noise models under the null (CURN) hypothesis. While the data are fixed, the models are uncertain and vary across posteriors, so we adopt a noise-marginalized approach [30, 29]. For each posterior-sample CURN universe, we compute the three statistics and evaluate their false-alarm probabilities using the corresponding generalized-$\chi^{2}$ distribution [34, 33]. We then form a histogram of the resulting $p$-values.
We first discuss the NPMV statistic, whose posterior $p$-value histogram is shown in Fig. 3. The mean/median false-alarm probabilities are $p=5.5\times 10^{-4}$$p=1.1\times 10^{-4}$ for free $\gamma_{gw}$ and $p=3.7\times 10^{-4}$$p=8.0\times 10^{-5}$ for fixed $\gamma_{gw}=13/3$. Respectively, these correspond to $3.3\sigma$$3.7\sigma$ and $3.4\sigma$$3.8\sigma$ equivalent Gaussian significances. The most likely range is shown in Fig. 3 as the band that contains 68% of the posterior points about the median. These correspond to significances in the range from $3.1\sigma$ to $4.3\sigma$ for free $\gamma_{gw}$ and $3.3\sigma$ to $4.3\sigma$ for $\gamma_{gw}=13/3$.
@@
OS $\sigma-$units: 4.3 4.8 (4.3, 5.4) (3.9, 5.9)
$\log_{10}p$: 5.2 6.2 (5.1, 7.4) (4.3, 8.7)
NP $\sigma-$units: 3.3 3.8 (3.2, 4.3) (2.7, 4.9)
$\log_{10}p$: 3.3 4.1 (3.2, 5.1) (2.4, 6.3)
NPMV $\sigma-$units: 3.3 3.7 (3.1, 4.3) (2.7, 4.9)
$\log_{10}p$: 3.3 4.0 (3.1, 5.1) (2.4, 6.3)
@@
The corresponding histograms for the NP and OS false-alarm probabilities $p$ are similar (see SM Sec. II). We list key results for all three statistics in Table 3. None has a mean false-alarm probability $p$ small enough to exceed the $5\sigma$ detection threshold [12]. However, note that our analysis yields a traditional OS detection significance of $\approx 4.3\sigma$, which is higher than the $\approx 3.2\sigma$ significance reported by NG [35]. This suggests that the “direct data combination” method is working as would be expected in the presence of an SGWB.
Another way to quantify detection confidence [12] is via the Bayes factor ${\mathrm{BF}}=E_{\rm{HD}}/E_{\rm{CURN}}$, the ratio of HD to CURN model evidences. Here, the evidence $E$ is the likelihood marginalized over the prior [59]; $E$ has units $({\rm time})^{-N}$ where $N=1{\kern 0.83298pt}090{\kern 0.83298pt}206$ is the number of TOAs (see SM Sec. I Table 4). Evaluating these integrals is challenging because the parameter space has $4\times 121+2=486$ dimensions for $\gamma_{gw}$ free ($485$ for $\gamma_{gw}=13/3$). To compute the evidences we use generalized stepping-stone sampling [60, 61, 62], which broadens the sharply peaked likelihood using $K=8$ chains (K1=7K-1=7 steps) while restricting the integration region. This gives $\ln(E_{\rm{HD}}\,\rm{s}^{N})=12{\kern 0.83298pt}500{\kern 0.83298pt}504.39\pm 0.10$ and $\ln(E_{\rm{CURN}}\,\rm{s}^{N})=12{\kern 0.83298pt}500{\kern 0.83298pt}494.21\pm 0.09$, implying $\ln({\mathrm{BF}})=10.18\pm 0.13$ and ${\mathrm{BF}}=26{\kern 0.83298pt}000\pm 3{\kern 0.83298pt}000$. This is two orders-of-magnitude larger than the Bayes factor ${\mathrm{BF}}\approx 200$ reported by NG [9]. For a real Gaussian random variable, the corresponding significance [29, Eq. (8)] of $\sqrt{2\ln({\mathrm{BF}})}\,\sigma\approx 4.5\sigma$ is below the threshold for confident detection.
@@
We test if the inter-pulsar correlations match those expected from an SGWB by estimating the angular correlation pattern from the data [12]. The mean correlation $\Gamma_{ab}$ between pulsars $a$ and $b$ arising from an SGWB should follow the HD prediction [26, 27, 2]
@@
Our HD reconstruction uses the NG [9] choice of 15 angular bins. The $121\times 120/2=7{\kern 0.83298pt}260$ pulsar pairs are divided equally: each bin has 484 pairs. For each posterior sample, the correlations of the pulsar pairs are weighted as in [64], giving $20{\kern 0.83298pt}480$ mean correlations $\mu$ and their variances $\sigma^{2}_{\mu}$ per bin. For each bin, we then construct and average the $20{\kern 0.83298pt}480$ corresponding Gaussian probability distributions. The resulting 15 (posterior estimators of the) probability distributions of the correlation are sharply peaked. Fig. 4 shows their means (close to medians) and standard deviations. A traditional $\chi^{2}$ goodness-of-fit test shows that the reconstruction matches the HD curve well: the reduced $\chi^{2}$ of $0.74$ is consistent with expected statistical fluctuations for 15 degrees of freedom
Conclusion — We search for an SGWB using public data from five PTAs. In effect, we repeat an established analysis, but with a larger dataset. This is enabled by the physically motivated step of adopting consistent astrophysical parameters and noise models for all pulsars.
@@
Our results provide the strongest evidence to date that an SGWB contributes to observed pulsar pulse-arrival-time fluctuations. However, the statistical significance—–assessed using multiple frequentist and Bayesian methods—–remains too low to claim a detection.
@@
Using existing analysis pipelines enables straightforward comparison with individual PTA results, with which our findings are consistent, but it has limitations. For example, these pipelines assume that for a stationary process, the Fourier amplitudes in different frequency bins are statistically independent. Although this assumption is not strictly valid [63, see Conclusion], its impact may be negligible or correctable [45]. Likewise, our HD-curve reconstruction is suboptimal because implementation of an optimal method [63] is still in progress [46]. Another limitation is the lack of a hierarchical pulsar-noise model [65, 66], which could shift the inferred SGWB amplitude and spectral slope [43].
@@
For its upcoming DR3 release, the IPTA [23] is reanalyzing the five PTA plus CPTA dataset to make new .par and .tim files. If the primary outcome is adopting common astrophysical models for the pulsars, then—per the arguments above—we expect an SGWB search of the DR3 five PTA data to yield results consistent with ours. But if reanalysis reveals that some data are incorrect and must be modified or discarded, the results could differ. Thus, the reliability of our conclusions depends on the correctness of the five existing PTA analyses.
@@
We deliberately kept this analysis straightforward and did not “iterate” it. This departs from standard pulsar-astronomy practice, where models (e.g., for pulsar timing) are iteratively revised. Here, we avoid iterative refinement and tuning of the full PTA search pipeline. Past experience with weak-signal detection in LIGO, as well as theoretical studies for PTAs [11], suggests that such iteration can be misleading.
@@
Because they are not public, we could not include CPTA data [8] in our analysis. These are from the Five-hundred-meter Aperture Spherical Telescope (FAST) in Guizhou, China, which is the world’s most sensitive radio telescope. Including these high-quality FAST data in future analyses would be interesting and informative.
Notes: (1) Our full analysis code will be released publicly upon publication. (2) After this work was circulated, a different data-combination method was proposed [67]. Unlike ours, this does not use a physically consistent timing model for a given pulsar. See also SM Sec. I.
@@
Acknowledgments – We thank R. van Haasteren for many useful conversations, encouragement on this project, and comments on manuscript drafts. We also thank S. Babak, S. Chen, C. Clark, O. Dodge, B. Goncharov, M. Kramer, P.M. Meyers, M.A. Papa, J.D. Romano, and M. Vallisneri for helpful feedback.
Supplementary Material
Here, we provide additional plots and tables as supporting material.
I Effects discarding “outlier backend” TOAs
In this work, we analyze the full five PTA dataset, which has $1{\kern 0.83298pt}090{\kern 0.83298pt}206$ TOAs. We were curious to see if the results would change if we discarded “outlier” data. By this, we mean data for which the single pulsar analysis gives WN parameters differing significantly from those naively implied by the .tim file contents. We flag such outliers via $\mathrm{EFAC}$}, whose nominal value is (close to) unity, as shown in Fig. 5.
Figure 5: Histograms showing how many backends and TOAs have a given value of $\mathrm{EFAC}$}.
To see the effects of discarding outliers, we drop TOAs from backends where $\mathrm{EFAC}$} lies outside the range $[1/Q,Q]$ for $Q=10,4,3$ and $2$. The number of backends/TOAs and fraction of data eliminated by these cuts are shown in Table 4.
Table 4: Summary of EFAC cuts and their impact on backends and TOAs.
EFAC Backends Backends TOAs TOAs % TOAs
range discarded retained discarded retained discarded
$0,\infty)$ 0 976 0 $1{\kern 0.83298pt}090{\kern 0.83298pt}206$ 0
$[0.10,10]$ 7 PPTA 969 $1{\kern 0.83298pt}915$ $1{\kern 0.83298pt}088{\kern 0.83298pt}291$ 0.18
$[0.25,4]$ 6 InPTA, 1 NG, 19 PPTA 960 $8{\kern 0.83298pt}226$ $1{\kern 0.83298pt}081{\kern 0.83298pt}980$ 0.75
$[0.33,3]$ 12 InPTA, 2 NG, 12 PPTA 950 $17{\kern 0.83298pt}002$ $1{\kern 0.83298pt}073{\kern 0.83298pt}204$ 1.56
$[0.50,2]$ 1 EPTA, 24 InPTA, 4 NG, 18 PPTA 929 $21{\kern 0.83298pt}340$ $1{\kern 0.83298pt}068{\kern 0.83298pt}866$ 1.96
The impact of dropping these “outlier” TOAs is negligible, as can be seen from the $A_{gw}$ and $\gamma_{gw}$ posteriors shown in Fig. 6. These should be compared with Fig. 2 in the paper. The tighter data cuts do not significantly affect the posteriors or other conclusions. We conclude that our “all are welcome” approach to data inclusion has not significantly skewed or affected our results, but note that discarding the “outlier” data systematically increases the inferred amplitude of the SGWB.
Figure 6: We illustrate the effects of using different $\mathrm{EFAC}$} thresholds to eliminate data from the most poorly modeled backends. The black line shows the CURN posterior from Fig. 2 in the paper. The other curves are constructed after dropping TOAs as shown in Table 4, and demonstrate that the CURN posteriors are not significantly affected by the data cuts. Note: each curve shows $20{\kern 0.83298pt}480$ posterior samples.
A recently posted paper [67] comments about “anomalous WN parameters” in a preliminary draft of this work. These arose because that earlier analysis contained phase disconnections between some TOAs observed with different backends. This was also the reason that 5 pulsars were given “special” treatment there. These phase disconnection issues were fixed by adding rotation counts to the input files.
II False-alarm probability $p$-value distributions
In the paper, Figure 3 shows a histogram of the posterior values of the false-alarm probability for the NPMV detection statistic. Here, in Figures 7 and 8, we provide the corresponding posterior $p$-value histograms for the Neyman-Pearson (NP) statistic and the traditional “optimal statistic” (OS). The means, medians, and widths of these distributions are reported in Table 3 of the paper.
Figure 7: Histograms of NP detection statistic $p$-values for the $20{\kern 0.83298pt}480$ CURN-hypothesis posterior samples. The left panel has $\gamma_{gw}$ free, and the right has $\gamma_{gw}=13/3$. The upper horizontal axis is one-sided-Gaussian equivalent significance.
Figure 8: Histograms of the traditional optimal detection statistic (OS) $p$-values for the $20{\kern 0.83298pt}480$ CURN-hypothesis posterior samples. The left panel has $\gamma_{gw}$ free, and the right has $\gamma_{gw}=13/3$. The upper horizontal axis is one-sided-Gaussian equivalent significance.
III Data Analysis Workflow
Figure 9: A graphical illustration of the data analysis workflow. The input data are .par and .tim files from the five PTAs. From these, we produce feather files ($F$), a white noise dictionary (WND), and posterior samples (PS) which are used for further analysis. The circled numbers one to five order the analysis steps.
Our data analysis workflow is illustrated in Fig. 9. It begins at the top left with .par and .tim files coming from the five PTAs’ public data releases. The .par files contain parameters such as the pulsar sky position and spin frequency, and the .tim files contain the TOAs.
The PTAs employ pulsar timing programs to prepare these files. These programs read and write the files, and are used to fit astrophysical parameters to the data, by minimizing the timing residuals. NG uses the PINT for this purpose, whereas the other four PTAs use Tempo2.
The direct data combination method adopts a common astrophysical model for each pulsar. In practice, we do this by modifying the .par files, as described in detail in the paper. To analyze the data for a given pulsar, we use a single system of units. For example, if some of the .par files for a given pulsar have TCB time coordinates, and others have TDB time coordinates, then we convert all of them to TDB. Quantities that contain units of time are scaled in value by this conversion. For example, pulsar rotation frequencies expressed in TDB are larger than in TCB by about 16 parts per billion. Alternatively, if all of the .par files use TDB or TCB exclusively, then time-coordinate-dependent quantities are not modified.
The situation is simpler for the sky position units. Sky positions can be specified using right ascension (RA) and declination (DEC) or in ecliptic latitude and longitude, and the angles can be given in hour-angles, degrees, or radians. For constructing a common timing design matrix $\mathbf{M}$}, we need to ensure that only one combination of these units is employed for a given pulsar. This is accomplished when the sky position is copied from the reference to the target .par files.
For .par files that specify the ELL1H binary model, there is an additional complication. In PINT, if the Shapiro delay parameter $H3$ is not specified, then it defaults to zero. However, until a recent revision, this was incompatible with Tempo2, which did not set a default value, but instead read a random value from uninitialized memory. Our solution: if there is no $H3$ specified for the ELL1H model, then we set binary model to ELL1 model, which is equivalent to ELL1H with $H3=0$.
Note that while PINT and Tempo2 internally use slightly different values for physical constants such as the solar mass, this makes no difference in our analysis.
The modified .par files and the .tim files (with pulse counts added) are then passed to Enterprise. This computes the timing residuals and design matrix $\mathbf{M}$} for each pulsar, which are stored in a single Feather file for each pulsar. The timing residuals are computed using PINT for the NG TOAs, and Tempo2 for the other TOAs. These timing residuals are what remain after subtracting an identical deterministic model for pulse arrival time from all TOAs of all PTAs. The Feather files, one per pulsar, are the inputs for all remaining analysis, as shown in Fig. 9. The .par and .tim files are no longer needed.
Table 5: Parameter name equivalences between PINT and Tempo2.
Physical meaning PINT Tempo2
Derivative of projected semi-major axis A1DOT XDOT
Eccentricity ECC E
The first time derivative of eccentricity EDOT ECCDOT
Shapiro delay parameter STIGMA STIG, VARSTIGMA
We perform a “sanity test”, to check that for a given pulsar, the separate Enterprise objects produced for each PTA have identical astrophysical model parameters. Although these parameters were identical in the .par files read by Enterprise, if the resulting objects are queried, the names in those produced with PINT are different than in those objects produced by Tempo2. The correspondence is shown in Table 5. So, after checking that the parameters have identical values and equivalent names, the Enterprise objects are modified to recognize all of the names. This ensures that the design matrix $\mathbf{M}$} contains the correct derivatives in any given column.